OpenStudy (mateeldaa):

f(x)= x^2/(2+x) f'(x)=4x+x^2/((2+x)^2) How do I go about finding the second derivative?

OpenStudy (anonymous):

u know how to differentiate ?

OpenStudy (mateeldaa):

we covered it in class, but I don't really understand it

OpenStudy (anonymous):

u can differentiate f(x) two times and u will get second derivative

OpenStudy (amistre64):

product rule tends to be simpler to assess [ab^(-1)]' = a'b^(-1) - a b' b^(-2)

OpenStudy (mateeldaa):

for f'(x)=\[\frac{ 4x+x ^{2} }{ (2+x)^{2} }\] this is what I got

OpenStudy (amistre64):

ugh ... wolf says its bad ... accursed algebra!!

OpenStudy (amistre64):

quotient rule: bt' - b't ------ b^2

OpenStudy (mateeldaa):

why can't math be easy??! haha

OpenStudy (amistre64):

\[\frac{ 4x+x ^{2} }{ (2+x)^{2} }\] \[\frac{ (2+x)^{2}(4x+x ^{2})'- (4x+x ^{2})[(2+x)^{2}]'}{ [(2+x)^{2} ]^2}\] \[\frac{ (2+x)^{2}(4+2x)- 2(4x+x ^{2})(2+x)}{ (2+x)^{4}}\] \[\frac{ (2+x)(4+2x)- 2(4x+x ^{2})}{ (2+x)^{3}}\] and simplify

OpenStudy (haseeb96):

@mateeldaa do u still need any help?

OpenStudy (mateeldaa):

thank y'all so much!!